Quinine has been used as an anti-malarial drug for many years. Its aqueous solution in certain acids has been used as a standard in the determination of fluorescence quantum yield of other fluorophores. The structure of quinine is shown below. Quinine in a 1.664-g antimalarial tablet was dissolved in sufficient 0.10 M HCl to give 500 mL of solution. A 10.00 mL aliquot was then diluted to 100.00 mL with the same acid. The fluorescence intensity for the diluted sample at 348 nm is 270 on an arbitrary scale. A standard 10.0 ?M quinine solution gave an intensity value of 180 at 348 nm when measured under conditions identical to those for the diluted sample. The blank solvent gave a reading of 0. There was no deviation from the linear relationship in fluorescence spectrometry. Nothing other than quinine was fluorescent in the tablet. ) Calculate the mass of quinine in the tablet (unite: mg).
Answer #1

concentration of quinine standard = 10 uM = 10 x 10^-6 M

intensity of standard = 180

molar absorptivity = 180/10 x 10^-6 = 1.8 x 10^7 M-1.cm-1

intensity of quinine diluted solution = 270

concentration of quinine in diluted solution = 270/1.8 x 10^7 = 1.5 x 10^-5 M

quinine in tablet = 1.5 x 10^-5 M x 100/10 = 1.5 x 10^-4 M

mass quinine in tablet = 1.5 x 10^-4 M x 0.5 L x 324.42 g/mol

                                    = 0.024 g = 24 mg

Using,